OpenStudy (anonymous):

Sonja and Dante are hiking up the side of a mountain. They reach the peak of 727 feet at the end of the day. If the mountain forms a 77° angle with the ground, what is the distance they hiked? Round your answer to the nearest foot. A. 168 feet B. 708 feet C. 746 feet D. 3232 feet

OpenStudy (dumbcow):



OpenStudy (anonymous):

I'm confused so it's A? @dumbcow

OpenStudy (anonymous):

I'm not good at these :/

OpenStudy (dumbcow):

i gave you the equation, solve for x for these problems remember SOH CAH TOA sin = opp/hyp cos = adj/hyp tan = opp/adj